# CoLo-CAM: Class Activation Mapping for Object Co-Localization in Weakly-Labeled Unconstrained Videos

### v4

@@ 3.1 Per-Frame Losses
a) Learning via pseudo-labels (PLs). It is commonly known that strong activations in a CAM are more likely to be a foreground, while low activations are assumed to be a background [87]. Following this, pixel-wise PLs for foreground and background regions are generated using the CAM ${\bm{C}_{t}}$ generated by the classifier ${g}$. ROIs are stochastically sampled instead of fixed ROIs [39]. This has been shown to be more effective and avoids overfitting [6]. Therefore, at each SGD step, a foreground pixel is sampled using a multinomial distribution over strong activations assuming that an object is local. Uniform distribution over low activations is used to sample background pixels assuming that background regions are evenly distributed in an image. The location of these two random pixels is encoded in ${\Omega^{\prime}_{t}}$. The partially pseudo-labeled mask for the sample ${\bm{X}_{t}}$ is denoted ${\bm{Y}_{t}}$, where ${\bm{Y}_{t}(p)\in\{0,1\}^{2}}$ at location $p$ with value 00{0} for background, 11{1}1 for foreground, and locations with unknown labels are encoded as unknown. At this stage, since the pre-trained classifier $f$ is frozen, its CAM ${\bm{C}_{t}}$ is fixed and does not change during the training of the decoder $g$, allowing for more stable sampling. The decoder output ${\bm{S}_{t}}$ is then aligned with ${\bm{Y}_{t}}$ using partial cross-entropy,
@@
where ${\bm{W}_{t}}$ is an affinity matrix in which ${\bm{W}[i,j]}$ captures the proximity between pair of pixels ${i,j}$ in the single frame ${\bm{X}_{t}}$, and their color similarity; ${r=0}$ is the index of the background, while ${r=1}$ is for the foreground. Since ${\bm{S}_{t}}$ has the same size as the input image, it is beneficial to minimize Eq.2, since pixels’ details is available. This allows to more accurately compare their color.
c) Absolute size constraint (ASC). Unbalanced activations is a common issue in CAMs [17, 8, 62]. Often, strong activations cover only small and most discriminative parts of an object, allowing the background to dominate. Alternatively, large parts of an image are activated as foreground [62]. To avoid such unbalance, global size prior over the CAM is employed. It consists of constraining the size of foreground and background regions to be as large as possible in a competitive way. To this end, Absolute Size Constraint (ASC) [7] on the decoder maps ${\bm{S}_{t}}$ is used. This is achieved without requiring any knowledge of the true object size or prior knowledge of which region is larger [58]. This prior is formulated as inequality constraints which are then solved via a standard log-barrier method [10],
$\sum\bm{S}^{r}_{t}\geq 0\;,\quad r\in\{0,1\}\;,$ (3)
where ${\psi(\bm{S}^{0}_{t})=\sum\bm{S}^{0}_{t},\psi(\bm{S}^{1}_{t})}$ represent the area, i.e., size, of the background (${r=0}$) and foreground (${r=1}$) regions, respectively. Using log-barrier function, we set,
$\mathcal{R}_{s}(\bm{S}_{t})=\sum_{r\in\{0,1\}}-\frac{1}{z}\log(\psi(\bm{S}^{r}_{t}))\;,$ (4)
where ${z>0}$ is a weight that is increased periodically; ${r=0}$ is for the background, while ${r=1}$ is for the foreground.
@@ 3.2 Multi-Frame Loss Based on Color Cues
Our new temporal loss extends over a sequence of frames. It builds a fully connected graph over all pixel locations across all frames. Using discriminative response over the decoder CAMs, and only the pairwise color similarity, a CRF energy loss [72] is set to be minimized. This amounts to achieve co-localization across multiple frames using color appearance. The emerged object localization at frame level via per-frame terms are propagated through all the frames, and ensured to be consistent. Regions that have been deemed to be foreground are maintained as foreground across all frames, and the same way for background. Such knowledge transfer allows reducing errors caused by noisy pseudo-labels at frame level leading to better localization. Our approach is depicted in Fig.2. Our loss is based on the assumption that locally, i.e. over the extent of few frames, objects and overall frames maintain the same color, which is a fair assumption in video processing. Unlike TCAM [5], we do not assume limited displacement of objects. This gives our method more flexibility to localize an object independently of its location.
@@
where ${\bm{W}}$ is the color-only similarity matrix between pixels of the concatenated image ${\bm{\bar{X}}}$. We note the main difference between CRF loss in Eq.2 which operates only on a single frame level, and accounts for both pairwise proximity and color similarity between pixels. On the other hand, our temporal CRF loss in Eq.5 operates on multiple frames, and acts only on pairwise color similarity across all frames without accounting for proximity between pixels or frames. This allows the object to be anywhere in the sequence or within a frame yielding more localization flexibility compared to TCAM method [5]. In our implementation, we use a Gaussian kernel to capture color similarities [41]. The kernel is implemented via the permutohedral lattice [1] for fast computation. We refer to the Eq.5 as CoLoc term.
@@ 3.3 Total Training Loss
$\min_{\bm{\theta}}\quad$ $\sum_{p\in\Omega^{\prime}_{t}}\bm{H}_{p}(\bm{Y}_{t},\bm{S}_{t})+\lambda\;\mathcal{R}(\bm{S}_{t},\bm{X}_{t})+\mathcal{R}_{s}(\bm{S}_{t})$ (6)
@@
After being trained using Eq. 6, our model is applied on single frames using a forward pass for inference (Fig.1). This allows for fast inference and parallel computation over a video stream. Using standard procedures [5, 17], a localization bounding box is extracted from the foreground CAM of the decoder ${\bm{S}^{1}_{t}}$. Our method yields a single model with parameters ${\{\bm{\theta^{\prime}},\bm{\theta}\}}$ performing both tasks, classification and localization, while handling multiple classes at once.
@@
4 Results and Discussion
4.1 Experimental Methodology
YouTube-Object v2.2 (YTOv2.2[36]: This is a large dataset built on YTOv1 dataset with large and challenging test set. It is composed of more frames, 722,040 in total. The dataset was divided by the authors into 106 videos for training, and 49 videos for test. Following [5], we sampled 3 random videos per class from the train set to build a validation set. The test set has more samples compared to YTOv1. It is composed of a total of 2,667 bounding box making it a much more challenging dataset.
Implementation Details. For all our experiments, we trained for 10 epochs with a mini-batch size of 32. ResNet50 [32] was used as a backbone. Further ablations are done with VGG16 [66], InceptionV3 [69]. Unless mentioned differently, all results are obtained using ResNet50 backbone. Images are resized to ${256\times 256}$, and randomly cropped patches of size ${224\times 224}$ are used for training. The temporal dependency ${n}$ in Eq.5 is set through the validation set from ${n\in\{2,\cdots,18\}}$. The hyper-parameter ${\lambda_{c}}$ in Eq.6 is set from ${\{1,\cdots,16\}}$ with the adaptive setup. In Eq.5, frames are concatenated horizontally to build ${\bm{\bar{X}}}$. We set the CRF hyper-parameter $\lambda$ in Eq.6 to ${2e^{-9}}$ – the same value as in [72]. Its color and spatial kernel bandwidth are set to 15 and 100, respectively, following [72]. The same color bandwidth is used for Eq.5. The initial value of the log-barrier coefficient, $z$ in Eq.4, is set to $1$ and then increases by a factor of 1.011.01{1.01}1.01 in each epoch with a maximum value of $10$ as in [4, 37]. The best learning rate was selected within ${\{0.1,0.01,0.001\}}$. The classifier $g$ was pre-trained on single frames on the train set.
Baseline Methods. For comparison, publicly available results were considered. We compared our proposed approach with different WSVOL methods [19, 29, 35, 42, 57, 59, 60, 73, 74], POD [38], SPFTN [81], and FPPVOS [75]. We also considered the performance of several CAM-based methods reported in [5]: CAM [87], GradCAM [64], GradCam++ [14], Smooth-GradCAM++ [54], XGradCAM [26], and LayerCAM [34], which were all trained on single frames, i.e., no temporal dependency was considered. We compare our results as well to CLIP-ES [45] using true class as input, and without background list or CAM-refinement for a fair comparison. The LayerCAM method [34] was employed to generate classifier CAMs ${\bm{C}_{t}}$ used to create pseudo-labels (${\bm{Y}_{t}}$) for our method (Eq.1). Note that different CAM-based methods can be integrated with our approach.
Evaluation Metric. Localization accuracy was evaluated using the CorLoc metric [5, 21]. It describes the percentage of predicted bounding boxes that have an Intersection Over Union (IoU) between prediction and ground truth greater than half (IoU ${>50\%}$). In addition, the inference time per frame of size ${224\times 224}$ is reported over GPU (NVIDIA Tesla P100) and CPU (Intel(R) Xeon(R), 48 cores). The same hardware is used to report the computation time of our multi-frame loss (Eq.5).
@@ 4.2 Comparison with the State-of-the-Art
The results in Tab.1 show that our CoLo-CAM method outperforms other methods by a large margin, more so on YTOv1 compared to YTOv2.2. This highlights the difficulty of YTOv2.2. In terms of per-class performance, our method is competitive in most classes. In particular, we observe that our method achieved a large improvement over the challenging class ’Train’. In general, our method and CAM methods are still behind on the class ’Horse’ and ’Aero’ compared to the SPFTN method [81] which relies on optical flow to generate proposals. Both classes present different challenges. The former class shows with dense multi-instances, while the latter, i.e., ’Aero’, appears in complex scenes (airport), often with very large size, and occlusion. In term of inference time, our method is relatively fast which is similar to TCAM [5] and F-CAM [8].
@@
4.3 Ablation Studies
The following experiments provide insight on how localization performance is affected by loss terms, backbone, sampling strategy. Computation time of our multi-frame loss is reported as well.
@@ 4.3 Ablation Studies
4.4 Analysis of CLIP-ES [45] Localization Performance
[…]
@@ 4.5 Visual Results
Figure 5: Typical challenges of our method. Top row: Dense and overlapped instances lead to fused localization of instances. Bottom row: Object mis-localization. Classifier CAM does not activate over the right object. Our trained decoder using these CAMs is unable to correct this large error. Bounding boxes: ground truth (green), prediction (red).
@@
Demonstrative videos provided with the code highlight additional challenges that are not visible in still images, in particular:
@@ 5 Conclusion
This paper introduces a new CAM-based approach for WSVOL. Using pixel-color cue, our proposed CoLo-CAM method constrains CAM responses to be similar over visually similar pixels across all frames, assuming that an object maintains a similar color over adjacent frames. Multi-frame optimization is achieved by constructing a fully connected graph between all pairs of pixels across all the considered frames. CoLo-CAM relies on the pixel probability of CAMs as unary potentials, and the color difference between every two pixels as the pairwise potentials. This allows for interconnection among all the pixels and performing explicit co-localization among frames. Our total training loss comprises per-frame and multi-frame terms optimized simultaneously via standard Stochastic Gradient Descent (SGD). Empirical experiments showed that our produced CAMs become more discriminative. This translates into several advantageous properties: sharp, more complete, and less noisy activations localized more precisely around objects. Moreover, localizing small/large objects becomes more accurate. Additionally, our method is more robust to training with long time-dependency and leads to new state-of-the-art localization performance.

### v5

@@ 3.1 Per-Frame Losses
a) Learning via pseudo-labels (PLs). It is commonly known that strong activations in a CAM are more likely to be a foreground, while low activations are assumed to be a background [99]. Following this, pixel-wise PLs for foreground and background regions are generated using the CAM ${\bm{C}_{t}}$ produced by the classifier ${g}$. ROIs are stochastically sampled instead of fixed ROIs [37]. This has been shown to be more effective and avoids overfitting [4]. Therefore, at each SGD step, a foreground pixel is sampled using a multinomial distribution over strong activations assuming that an object is local. Uniform distribution over low activations is used to sample background pixels assuming that background regions are evenly distributed in an image. The location of these two random pixels is encoded in ${\Omega^{\prime}_{t}}$. The partially pseudo-labeled mask for the sample ${\bm{X}_{t}}$ is denoted ${\bm{Y}_{t}}$, where ${\bm{Y}_{t}(p)\in\{0,1\}^{2}}$ at location $p$ with value 00{0} for background, 11{1}1 for foreground, and locations with unknown labels are encoded as unknown. At this stage, since the pre-trained classifier $f$ is frozen, its CAM ${\bm{C}_{t}}$ is fixed and does not change during the training of the decoder $f$, allowing for more stable sampling. The decoder output ${\bm{S}_{t}}$ is then aligned with ${\bm{Y}_{t}}$ using partial cross-entropy,
@@
where ${\bm{W}_{t}}$ is an affinity matrix in which ${\bm{W}[i,j]}$ captures the proximity between pair of pixels ${i,j}$ in the single frame ${\bm{X}_{t}}$, and their color similarity; ${r=0}$ is the index of the background, while ${r=1}$ is for the foreground. Since ${\bm{S}_{t}}$ has the same size as the input image, it is beneficial to minimize Eq.2, since pixels’ details is available, allowing for more accurate pixel-wise comparison.
c) Absolute size constraint (ASC). Unbalanced activations is a common issue in CAMs [12, 63, 7]. Often, strong activations cover only small and most discriminative parts of an object, allowing the background to dominate. Alternatively, large parts of an image are activated as foreground [63]. To avoid such unbalance, global size prior over the CAM is employed. It consists of constraining the size of foreground and background regions to be as large as possible in a competitive way. To this end, Absolute Size Constraint (ASC) [6] on the decoder maps ${\bm{S}_{t}}$ is employed. This is achieved without requiring any knowledge of the true object size or prior knowledge of which region is larger. In particular, ASC allows the foreground and background region to compete by pushing their area to be as large as possible, effectively preventing unbalanced activation. This prior is formulated as inequality constraints which are then solved via a standard log-barrier method by minimizing the following loss,
$\mathcal{R}_{s}(\bm{S}_{t})=-\frac{1}{z}\left[\log(\psi(\bm{S}^{0}_{t}))+\log(\psi(\bm{S}^{1}_{t}))\right]\;,$ (3)
where ${\psi(\bm{S}^{0}_{t})=\sum\bm{S}^{0}_{t}}$ represents the area, i.e., size, of the background, and ${\psi(\bm{S}^{0}_{t})}$ is the foreground area; ${z>0}$ is a weight that is increased periodically to strengthen the log-barrier loss.
@@ 3.2 Multi-Frame Loss Based on Color Cues
Our new temporal loss extends over a sequence of frames. It builds a fully connected graph over all pixel locations across all frames. Using discriminative response over the decoder’s CAMs, and only the pairwise color similarity, a CRF energy loss [76] is set to be minimized. This amounts to achieve co-localization across multiple frames using color appearance. The emerged object localization at frame level via per-frame terms are propagated through all the frames, and ensured to be consistent. Regions that have been deemed to be foreground are maintained as foreground across all frames, and the same way for background. Such knowledge transfer allows reducing errors caused by noisy pseudo-labels at frame level leading to better localization. Our approach is depicted in Fig.1. Our loss is based on the assumption that locally, i.e. over the extent of few frames, objects and overall frames maintain the same color, which is a fair assumption in video processing. Unlike TCAM [3], we do not assume limited displacement of objects. This gives our method more flexibility to localize an object independently of its location.
@@
where ${\bm{W}}$ is the color-only similarity matrix between pixels of the concatenated image ${\bm{\bar{X}}}$. We note the main difference between CRF loss in Eq.2 which operates only on a single frame level, and accounts for both pairwise proximity and color similarity between pixels. On the other hand, our temporal CRF loss in Eq.4 operates on multiple frames, and acts only on pairwise color similarity across all frames without accounting for proximity between pixels or frames. This allows the object to be anywhere in the sequence or within a frame yielding more localization flexibility compared to TCAM method [3]. In our implementation, we use a Gaussian kernel to capture color similarities. The kernel is implemented via the permutohedral lattice for fast computation. We refer to the Eq.4 as CoLoc term.
@@ 3.3 Total Training Loss
$\min_{\bm{\theta}}$ $\sum_{p\in\Omega^{\prime}_{t}}\bm{H}_{p}(\bm{Y}_{t},\bm{S}_{t})+\lambda\;\mathcal{R}(\bm{S}_{t},\bm{X}_{t})+\mathcal{R}_{s}(\bm{S}_{t})$ (5)
@@
After being trained using Eq. 5, our model is applied over single frames using a forward pass for inference (Fig.6). This allows for fast inference and parallel computation over a video stream. Using standard procedures [3, 12], a localization bounding box is extracted from the foreground CAM of the decoder ${\bm{S}^{1}_{t}}$. Our method yields a single model with parameters ${\{\bm{\theta^{\prime}},\bm{\theta}\}}$ performing both tasks, classification and localization, while handling multiple classes at once.
4 Results and Discussion
4.1 Experimental Methodology
YouTube-Object v2.2 (YTOv2.2[34]: This is a large dataset built on YTOv1 dataset with large and challenging test set. It is composed of more frames, 722,040 in total. The dataset was divided by the authors into 106 videos for training, and 49 videos for test. Following [3], we sampled 3 random videos per class from the train set to build a validation set for model selection via early stopping. The test set has more samples compared to YTOv1. It is composed of a total of 2,667 bounding box making it a much more challenging dataset. We note that both datasets came already divided into train and test set by their authors. We use the same splits. Only the validation set used for model selection is selected randomly from the provided train set, as done in previous works.
Implementation Details. For all our experiments, we trained for 10 epochs with a mini-batch size of 32. ResNet50 [29] was used as a backbone. Further ablations are done with VGG16 [68], InceptionV3 [73]. Unless mentioned differently, all results are obtained using ResNet50 backbone. Images are resized to ${256\times 256}$, and randomly cropped patches of size ${224\times 224}$ are used for training. The temporal dependency ${n}$ in Eq.4 is set through the validation set from ${n\in\{2,\cdots,18\}}$. The hyper-parameter ${\lambda_{c}}$ in Eq.5 is set from ${\{1,\cdots,16\}}$ with the adaptive setup. In Eq.4, frames are concatenated horizontally to build ${\bm{\bar{X}}}$. We set the CRF hyper-parameter $\lambda$ in Eq.5 to ${2e^{-9}}$ – the same value as in [76]. Its color and spatial kernel bandwidth are set to 15 and 100, respectively, following [76]. The same color bandwidth is used for Eq.4. The initial value of the log-barrier coefficient, $z$ in Eq.3, is set to $1$ and then increases by a factor of 1.011.01{1.01}1.01 in each epoch with a maximum value of $10$. The best learning rate was selected within ${\{0.1,0.01,0.001\}}$. The classifier $g$ was pre-trained on single frames on the train set.
Baseline Methods. For comparison, publicly available results were considered. We compared our proposed approach with different WSVOL methods [14, 27, 33, 38, 58, 60, 77, 78], POD [36], SPFTN [90], and FPPVOS [79]. We also considered the performance of several CAM-based methods reported in [3]: CAM [99], GradCAM [65], GradCam++ [10], Smooth-GradCAM++ [56], XGradCAM [23], and LayerCAM [32], which were all trained on single frames, i.e., no temporal dependency was considered. We compare our results as well to CLIP-ES [43] using true class as input, and without background list or CAM-refinement for a fair comparison. The LayerCAM method [32] was employed to generate classifier CAMs ${\bm{C}_{t}}$ used to create pseudo-labels (${\bm{Y}_{t}}$) for our method (Eq.1). Note that different CAM-based methods can be integrated with our approach.
Evaluation Metric. Localization accuracy was evaluated using the CorLoc metric [3]. It describes the percentage of predicted bounding boxes that have an Intersection Over Union (IoU) between prediction and ground truth greater than half (IoU ${>50\%}$). In addition, the inference time per frame of size ${224\times 224}$ is reported over GPU (NVIDIA Tesla P100) and CPU (Intel(R) Xeon(R), 48 cores). The same hardware is used to report the computation time of our multi-frame loss (Eq.4).
@@ 4.2 Comparison with the State-of-the-Art
The results in Tab.1 show that our CoLo-CAM method outperforms other methods by a large margin, more so on YTOv1 compared to YTOv2.2. This highlights the difficulty of YTOv2.2. In term of per-class performance, our method is competitive in most classes. In particular, we observe that our method achieved a large improvement over the challenging class ’Train’. In general, our method and CAM methods are still behind on the class ’Horse’ and ’Aero’ compared to the SPFTN method [90] which relies on optical flow to generate proposals. Both classes present different challenges. The former class shows with dense multi-instances, while the latter, i.e., ’Aero’, appears in complex scenes (airport), often with very large size, and occlusion. In term of inference time, our method is relatively fast which is similar to TCAM [3] and F-CAM [7].
4.3 Ablation Studies
The following experiments provide insight on how localization performance is affected by loss terms, backbone, sampling strategy. Computation time of our multi-frame loss is reported as well in addition to performance stability of our method.
@@ 4.3 Ablation Studies
We conducted furthermore analysis of the computational cost of our method during training and inference, while comparing to its baseline Layer-CAM method [32]. Table.4 shows the obtained results. Over training, the backward pass with all the losses and parameters update of our method have a high cost amounting to 0.84s for single SGD step while forward pass of ${n=64}$ reaches ${\sim 80}$. Its counterpart method achieves 0.06s. However, this training speed of our method remains practical. At inference time, our method is relatively faster than Layer-CAM since this last one requires gradient computation. For instance, our method can process image with a rate of 464 FPS compared to Layer-CAM rate of 128 FPS which over-loads the GPU memory.
Although, multiframe loss does not add a large computational overhead, it can be further sped up. One way to achieve this is to reduce the number of nodes in the fully connected graph. This can be done by considering a node as a small region such as an image superpixel or a patch. However, this required an adapted heuristic to describe the color of each region and reduce it to a single value similar to a pixel, along with the regions’ corresponding probabilities. Another way is to set the decoder to produce smaller localization maps than the full image size. This will effectively reduce the number of locations. In this case, the image can be simply interpolated to a small size.
Case/Approach Layer-CAM (w. ResNet50) Ours (w. ResNet50)
Train time 1 SGD step Batch size: 32 0.06s 0.84s
Inference time per-frame 19.99ms 10.11ms
Inference time: Frame rate (FPS) 128* 464
Total n. params. 23.5M 32.5M
N. learnable params. 23.5M 9.0M
N. FLOPs (GFLOPs) (1 frame) 38.23 55.49
N. MACs (GMACs) (1 frame) 19.07 27.70
Table 4: Comparison of computation time, number of parameters, number of FLOPs/MACs of basic model vs ours. Computations are performed on an NVIDIA A100. Frame size: ${224\times 224}$.
*: The frame rate (FPS) for Layer-CAM runs to an out of memory issue when going beyond 128 frames over a GPU with 40GB of memory. Processing 128 frames takes roughly 200ms. The FLOPs/MACs are computed using the library pypi.org/project/calflops (V0.3.2).
@@
g) Performance stability (Fig.2(c)). To further study the performance stability of our method, we conducted a new experiment by randomizing the effective train set since there are only very few labeled videos for validation set to be changed. Therefore, we keep the validation and test set fix. We only change the effective train set by discarding random shots for the main train set used so far. More specifically, we randomly removed $m$ shots per-class from the effective train set. We set ${m\in\{1,5,10,20,30,40,50,60,70,80,90,100\}}$. Note that each video has many shots (i.e., short duration clip). This leads to a new effective train set which is less favorable to our method since it has less data. However, the purpose of this experiment is to check stability of our method. Consequently, the obtained results are not directly comparable to our main result with full train set nor to previous works. This new experiment consists of repeating the process of discarding $m$ random shots 30 times over both datasets YTOv1 and YTOv2.2 using a fixed hyper-parameters set making the train set videos the only variable. Figure.3 shows the obtained results. We observe that in general, the results are fairly stable as they exhibit low standard deviation (Tab.7). This last one varies from 1.301.30{1.30}1.30 to 1.721.72{1.72}1.72 over YTOv1, and between 1.61.6{1.6}1.6 to 2.262.26{2.26}2.26 over YTOv2.2. Note that even when we drop 100 shots per class, our method still yields descent performance. However, as expected, reducing the size of effective train set leads to a drop in performance in general.
(a) Performance stability: YTOv1
(b) Performance stability: YTOv2.2
Figure 3: Stability of CoLo-CAM localization performance (CorLoc) on YTOv1 and YTOv2.2 test set. Results are obtained when randomizing the effective training set by removing $m$ random shots from the train set. This process is repeated 30 times.
$m$ / Dataset YTOv1 YTOv2.2
1 $78.68\pm 1.72$ $71.84\pm 1.66$
5 $78.07\pm 1.55$ $72.01\pm 2.17$
10 $78.65\pm 1.33$ $71.35\pm 1.84$
20 $78.26\pm 1.63$ $71.43\pm 1.90$
30 $77.48\pm 1.30$ $71.69\pm 1.89$
40 $77.84\pm 1.52$ $71.31\pm 1.63$
50 $77.73\pm 1.45$ $70.30\pm 1.92$
60 $77.35\pm 1.86$ $69.59\pm 1.60$
70 $77.22\pm 1.72$ $69.81\pm 2.26$
80 $76.85\pm 1.50$ $69.63\pm 1.84$
90 $76.28\pm 1.57$ $69.02\pm 2.19$
100 $75.77\pm 1.68$ $69.81\pm 2.11$
[…]
4.4 Analysis of CLIP-ES [43] Localization Performance
[…]
@@ 4.4 Analysis of CLIP-ES [43] Localization Performance
Figure 4: Typical challenges of our method. Column 1-2: Dense and overlapped instances lead to fused localization of instances. Column 3-4: Object mis-localization. Classifier CAM does not activate over the right object. Our trained decoder using these CAMs is unable to correct this large error. Bounding boxes: ground truth (green), prediction (red).
@@ 4.5 Visual Results
Demonstrative videos provided in the code highlight additional challenges that are not visible in still images, in particular:
@@ 5 Conclusion
This paper proposes a new CAM-based approach for WSVOL by introducing co-localization into a CAM-based method. Using pixel-color cue, our proposed CoLo-CAM method constrains CAM responses to be similar over visually similar pixels across all frames, achieving col-localization. Our total training loss comprises per-frame and multi-frame terms optimized simultaneously via standard Stochastic Gradient Descent (SGD). Empirical experiments showed that our produced CAMs become more discriminative. This translates into several advantageous properties: sharp, more complete, and less noisy activations localized more precisely around objects. Moreover, localizing small/large objects becomes more accurate. Additionally, our method is more robust to training with long time-dependency and leads to new state-of-the-art localization performance.
Despite its improvements, our method still holds some limitations. Our pixel pseudo-labels are less accurate since they come from a pretrained classifier. In particular, a major issue is the set of frames without object of interest in a video. In WSVOL, it is typically assumed that the video label transfers to each frame. However, in unconstrained videos, not all frames hold the labeled object. Therefore, the label is mistakenly transferred into empty frames. This furthermore allows the classifier to suggest relevant ROIs in such frames leading to more errors in training. One possible way to alleviate this issue is to leverage the per-class probabilities of the classifier to assess whether an object is present or not in a frame. Such likelihood can be exploited to discard frames or weight loss terms over them.
Another issue is related to inference that is performed frame by by frame without leveraging temporal information in a video. The negative impact of this can be observed over a sequence of predictions which are characterized by localization inconsistencies. This is illustrated by large shift in localization where the bounding box (and the CAM activation), moves abruptly from one location at a frame to completely different location at the next frame. Prediction without accounting for temporal dependencies can easily lead to such results. Future works may consider improving the inference procedure by leveraging spatiotemporal information at the expense of inference time.
We note that our method can use either CNN- or transformer-based architecture. The only requirement is that the architecture is able to perform both tasks: classification and localization. A vision transformer-based (ViT) model [16] can be employed as an encoder. A standard classification head can be used to perform classification task. A new localization head is required to replace our decoder to perform localization task. An architecture similar to the one presented in [53] can be used in our work which is expected to yield better performance compared to CNN-based model as reported in their work. We leave this aspect for future work.